# Uplift modeling with continuous treatments: A predict-then-optimize approach

### v1

@@ 3.3 Optimization step
The cost matrix $\mathbf{C}\in\mathbb{R}^{N\times\delta}$ defines treatment costs, where $c_{i,d}$ is the treatment cost for entity $i$ at dose $s$, with $c_{i,0}=0$. The simplest cost function assigns a cost directly proportional to the dose level, meaning a dose of $0.2$ incurs a cost of $0.2$. This cost function is used throughout the remainder of the paper and experiments for simplicity. However, it is modular and can be easily replaced with alternative cost structures to suit different scenarios or requirements. The cost of policy $\pi$ for instance $i$ is:
@@
$V_{i}^{exp}$ $=\langle\pi(\hat{\tau}(\mathbf{X}_{i})),\hat{\tau}(\mathbf{X}_{i})\rangle b_{i},$ (3)
$V_{i}^{presc}$ $=\langle\pi(\hat{\tau}(\mathbf{X}_{i})),\tau(\mathbf{X}_{i})\rangle b_{i},$ (4)
$V_{i}^{opt}$ $=\langle\pi(\tau(\mathbf{X}_{i})),\tau(\mathbf{X}_{i})\rangle b_{i}.$ (5)
For $N$ entities with features $\mathbf{X}$ and a budget $B$, the policy $\pi_{B}:\mathbb{R^{+}}\times\mathcal{X}^{N}\rightarrow\{0,1\}^{(N\times\delta)}$ can be based on estimated or ground-truth CADRs. The expected optimal policy $\pi_{B}^{exp}$ (Eq. 6) maximizes value within the budget using estimated CADRs. The prescribed policy assigns treatments according to $\pi_{B}^{exp}$, but uses ground-truth CADRs. The ground-truth optimal policy $\pi_{B}^{opt}$ (Eq. 7), based solely on ground-truth CADRs, is the full-information benchmark.
$\pi_{B}^{exp}=\operatorname{argmax}\left\{\sum_{i=1}^{N}V_{i}^{exp}(\pi):\sum_{i=1}^{N}\Psi_{i}(\pi)\leq B\right\}$ (6)
$\pi_{B}^{opt}=\operatorname{argmax}\left\{\sum_{i=1}^{N}V_{i}^{opt}(\pi):\sum_{i=1}^{N}\Psi_{i}(\pi)\leq B\right\}$ (7)
@@ 4.1 Predictive model for CADR estimation
To accurately estimate CADRs $\hat{\mu}$ and their associated CADE estimations $\hat{\tau}$, not addressing confounding between $S$ and $\bm{X}$ in observational data can lead to detrimental inaccuracies or biases (Pearl, 2009). A variety of methods for disentangling treatment effects from confounding factors has been proposed in the literature (Johansson et al., 2016; Schwab et al., 2020; Bockel-Rickermann et al., 2023), and this is not a prime concern in this paper. Instead, we consider the predictive methods as off-the-shelf solutions, with also methods being applied in the experiments that do not apply a debiasing scheme. One may refer to, e.g., Bica et al. (2020), for assessments of the impact of confounding on outcomes.
Following standard practices in the causal inference literature, we make three standard assumptions for estimating potential outcomes from observational data (Imbens, 2000; Lechner, 2001).
@@
Ignorability: $\left\{Y(s)\mid s\in\mathcal{S}\right\}\Perp S\mid\mathbf{X}$. This means that conditional on an entity’s pre-treatment characteristics, the assigned treatment dose $S$ is independent of the potential outcomes.
@@
In this work, we compare four different learning methods to estimate $\hat{\mu}$: an S-learner with random forests as the base learner (S-Learner (rf)), an S-learner with a feedforward multi-layer perceptron (MLP) without any debiasing as the base learner (S-Learner (mlp)) (Künzel et al., 2019), DRNet (Schwab et al., 2020), and VCNet (Nie et al., 2021). Further details on the hyperparameters can be found in D.
@@ 4.2 ILP for the dose-allocation problem
Deciding on the optimal treatment dose for each entity, represented by policies $\pi^{exp}_{B}$ and $\pi^{opt}_{B}$, is equivalent to solving Equations (6-7). In the case when only binary doses can be applied, i.e., $S\in\{0,1\}$, this problem is an instance of the knapsack problem (Dantzig, 1957). A sensible heuristic would be to rank allocations from high to low uplift until the budget capacity is reached. This approach mirrors the calculation of a traditional Qini curve in a binary treatment setting. In our setting, however, treatments are not restricted to binary values. Any dose $S\in[0,1]$ is possible. To accommodate this, we define a function $V_{B}:\{0,1\}^{(N\times\delta)}\times[-1,1]^{(N\times\delta)}\times\mathbb{R}^{N}\rightarrow\mathbb{R}$. This function takes as input a policy $\pi_{B}\in\{0,1\}^{(N\times\delta)}$ (expressed as a matrix of decision variables), a matrix $\in[-1,1]^{(N\times\delta)}$ of CADE values (either estimated $\hat{\tau}$ or ground-truth $\tau$) for $N$ entities, and a benefit vector $\mathbf{b}\in\mathbb{R}^{N}$, returning a value $V_{B}\in\mathbb{R}$. To solve Equations (6-7), we propose the ILP formulation below (8). The argument $V_{i}(\pi)$ in the objective function is set to $V_{i}^{exp}$ for finding $\pi_{B}^{exp}$ and to $V_{i}^{opt}$ for $\pi_{B}^{opt}$. Although numerous business requirements exist, we focus on fairness as the main example of constraints throughout this work. Without loss of generalization, we consider one protected binary sensitive attribute $A$, such as gender, where $N_{\mathrm{A=0}}$ is the index set of entities where the protected attribute $A=0$, and $N_{\mathrm{A=1}}$ represents those with $A=1$. These two sets are mutually exclusive and collectively exhaustive, so that $|N_{\mathrm{A=0}}|+|N_{\mathrm{A=1}}|=|N|$. To address fairness in allocation and outcomes, we introduce two slack parameters: $\epsilon_{s}\in[0,1]$ for allocation fairness and $\epsilon_{\tau}\in[0,1]$ for outcome fairness. A slack parameter value of $\epsilon=0$ guarantees perfect fairness between the groups $N_{\mathrm{A=0}}$ and $N_{\mathrm{A=1}}$, while parameter $\epsilon=1$ removes this fairness constraint.
$\max_{\pi_{B}}$ $\sum_{i=1}^{N}V_{i}(\pi)$ (8) s.t. $\sum_{i=1}^{N}\Psi_{i}(\pi)\leq B,$ (budget constr.) $\sum_{i\in{N_{\mathrm{A=0}}}}\sum_{s\in{D}}\frac{\pi_{s}(X_{i})*s}{|N_{\mathrm{A=0}}|}\geq(1-\epsilon_{s})\sum_{i\in{N_{\mathrm{A=1}}}}\sum_{s\in{D}}\frac{\pi_{s}(X_{i})*s}{|N_{\mathrm{A=1}}|}$ (alloc. fairn. pt.1) $\sum_{i\in{N_{\mathrm{A=0}}}}\sum_{s\in{D}}\frac{\pi_{s}(X_{i})*s}{|N_{\mathrm{A=0}}|}\leq(1+\epsilon_{s})\sum_{i\in{N_{\mathrm{A=1}}}}\sum_{s\in{D}}\frac{\pi_{s}(X_{i})*s}{|N_{\mathrm{A=1}}|}$ (alloc. fairn. pt.2) $\sum_{i\in{N_{\mathrm{A=0}}}}\sum_{s\in{D}}\frac{\pi_{s}(X_{i})*\tau_{s}}{|N_{\mathrm{A=0}}|}\geq(1-\epsilon_{\tau})\sum_{i\in{N_{\mathrm{A=1}}}}\sum_{s\in{D}}\frac{\pi_{s}(X_{i})*\tau_{s}}{|N_{\mathrm{A=1}}|}$ (outc. fairn. pt.1) $\sum_{i\in{N_{\mathrm{A=0}}}}\sum_{s\in{D}}\frac{\pi_{s}(X_{i})*\tau_{s}}{|N_{\mathrm{A=0}}|}\leq(1+\epsilon_{\tau})\sum_{i\in{N_{\mathrm{A=1}}}}\sum_{s\in{D}}\frac{\pi_{s}(X_{i})*\tau_{s}}{|N_{\mathrm{A=1}}|}$ (outc. fairn. pt.2) $\sum_{s\in{D}}\pi_{s}\left(X_{i}\right)=1,\forall i\in\{1\ldots N\},$ (each entity 1 dose) $\pi_{s}\left(X_{i}\right)\in\{0,1\},\forall s\in D,\forall i\in\{1,\ldots,N\}$ (int. constr.)
@@
Analogous to $V_{i}(\pi)^{exp}$, $V_{i}(\pi)^{presc}$, $V_{i}(\pi)^{opt}$ (Equations (3-5), which are defined on instance-level), and Equations (6-7) in Section 3, we consider three versions of $V_{B}$ each differing in their dependence on ground-truth or estimated CADEs:
$V_{B}^{exp}=V\left(\pi^{exp}_{B}(\hat{\tau}),\hat{\tau},\mathbf{b}\right),$ (9)
$V_{B}^{presc}=V\left(\pi^{exp}_{B}(\hat{\tau}),\tau,\mathbf{b}\right),$ (10)
$V_{B}^{opt}=V\left(\pi^{opt}_{B}(\tau),\tau,\mathbf{b}\right).$ (11)
@@ 5.1 Data
Our experiments use a semi-synthetic approach based on a real dataset about infant health and development programs (IHDP) (Brooks-Gunn et al., 1992). This dataset originates from a randomized controlled trial and is frequently used as a benchmark for binary CATE estimation methods (Shalit et al., 2017; Guo et al., 2020). Since we focus on continuous treatments, we adopt a semi-synthetic approach where both doses and outcomes are artificially generated. For this purpose, we follow the established literature (Nie et al., 2021). Detailed information about the original dataset and the semi-synthetic data generation process can be found in B.

### v2

@@ 3.3 Optimization step
The cost matrix $\mathbf{C}\in\mathbb{R}^{N\times\delta}$ defines treatment costs, where $c_{i,d}$ is the treatment cost for entity $i$ at dose $s$, with $c_{i,0}=0$. For each entity $i,$ its row in the cost matrix is denoted by $\mathbf{C}_{i}\;=\;(c_{i,0},\,c_{i,1},\,\dots,\,c_{i,\delta-1})$, which lists all possible costs for that entity. For simplicity, we assume costs are directly proportional to dose levels (e.g., a dose of $0.2$ has a cost of $0.2$). This cost function is modular and can be easily replaced with alternative cost structures to suit different scenarios or requirements. The cost of policy $\pi$ for instance $i$ is:
@@
Due to the fundamental problem of causal inference, historical data provides only factual outcomes, not counterfactual outcomes. Thus, in realistic scenarios, we must estimate these counterfactual outcomes. In contrast, experimental settings can leverage semi-synthetic datasets where the ground truth is known because we control the data-generating process. For a detailed discussion on the fundamental problem of causal inference, we refer readers to Pearl (2009). We define three distinct assignment policy values, each depending on the availability and usage of ground-truth versus estimated CADR values.
The expected policy value (Eq. 6) uses estimated CADRs for both the assignment policy and evaluation. It represents the anticipated value achievable under realistic conditions without access to counterfactual ground-truth data. The prescribed policy value $V_{i}^{presc}$ (Eq. 7) employs estimated CADRs to determine the assignment policy but uses ground-truth CADRs for evaluation. Policies determined by maximizing $V_{i}^{exp}$ yield an observable policy value $V_{i}^{presc}$. Due to differences between estimated and true causal effects, $V_{i}^{presc}$ typically diverges from $V_{i}^{exp}$. The optimal policy value $V_{i}^{opt}$ (Eq. 8) serves as a benchmark and uses ground-truth CADRs for both policy determination and evaluation. This represents the maximum attainable policy value in a hypothetical setting where all counterfactual dose-response information is known precisely, a condition unattainable in practical applications.
Formally, these are defined as:
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}U_{i}^{exp}}$ ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}=\langle\pi(\hat{\tau}(\mathbf{X}_{i})),\hat{\tau}(\mathbf{X}_{i})\rangle,}$ (3)
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}U_{i}^{presc}}$ ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}=\langle\pi(\hat{\tau}(\mathbf{X}_{i})),\tau(\mathbf{X}_{i})\rangle,}$ (4)
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}U_{i}^{opt}}$ ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}=\langle\pi(\tau(\mathbf{X}_{i})),\tau(\mathbf{X}_{i})\rangle.}$ (5)
$V_{i}^{exp}$ $=U_{i}^{exp}b_{i}=\langle\pi(\hat{\tau}(\mathbf{X}_{i})),\hat{\tau}(\mathbf{X}_{i})\rangle b_{i},$ (6)
$V_{i}^{presc}$ $=U_{i}^{presc}b_{i}=\langle\pi(\hat{\tau}(\mathbf{X}_{i})),\tau(\mathbf{X}_{i})\rangle b_{i},$ (7)
$V_{i}^{opt}$ $=U_{i}^{opt}b_{i}=\langle\pi(\tau(\mathbf{X}_{i})),\tau(\mathbf{X}_{i})\rangle b_{i}.$ (8)
For $N$ entities with features $\mathbf{X}$ and a budget $B$, the policy $\stackrel{{\scriptstyle B}}{{\Pi}}:\mathbb{R^{+}}\times\mathcal{X}^{N}\rightarrow\{0,1\}^{(N\times\delta)}$ can be based on estimated or ground-truth CADRs. The expected optimal policy $\stackrel{{\scriptstyle B}}{{\Pi}}^{exp}$ (Eq. 9) maximizes value within the budget using estimated CADRs. The prescribed policy assigns treatments according to $\stackrel{{\scriptstyle B}}{{\Pi}}^{exp}$, but uses ground-truth CADRs. The ground-truth optimal policy $\stackrel{{\scriptstyle B}}{{\Pi}}^{opt}$ (Eq. 10), based solely on ground-truth CADRs, is the full-information benchmark.
$\stackrel{{\scriptstyle B}}{{\Pi}}^{exp}=\operatorname{argmax}\left\{\sum_{i=1}^{N}V_{i}^{exp}(\pi):\sum_{i=1}^{N}\Psi_{i}(\pi)\leq B\right\}$ (9)
$\stackrel{{\scriptstyle B}}{{\Pi}}^{opt}=\operatorname{argmax}\left\{\sum_{i=1}^{N}V_{i}^{opt}(\pi):\sum_{i=1}^{N}\Psi_{i}(\pi)\leq B\right\}$ (10)
The implementation of this optimization formulation is discussed in Section 4.2.
@@ 4.1 Predictive model for CADR estimation
To accurately estimate CADRs $\hat{\mu}$ and their associated CADE estimations $\hat{\tau}$, not addressing confounding between $S$ and $\bm{X}$ in observational data can lead to detrimental inaccuracies or biases (Pearl, 2009). A variety of methods for disentangling treatment effects from confounding factors have been proposed in the literature (Johansson et al., 2016; Schwab et al., 2020; Bockel-Rickermann et al., 2023), and this is not a prime concern in this paper. Instead, we consider the predictive methods as off-the-shelf solutions, with methods that do not apply a debiasing scheme also being applied in the experiments. One may refer to, e.g., Bica et al. (2020), for assessments of the impact of confounding on outcomes.
Following standard practices in causal inference, we make three assumptions for estimating potential outcomes from observational data: Consistency, Ignorability, and Overlap (Imbens, 2000; Lechner, 2001). These assumptions ensure the identifiability of the CADE function $\tau_{s}(\mathbf{x})$, i.e., they result in Equation 29 (see proof in B).
@@
Ignorability: $\left\{Y(s)\mid s\in\mathcal{S}\right\}\Perp S\mid\mathbf{X}$. This means that, conditional on an entity’s pre-treatment characteristics, the assigned treatment dose $S$ is independent of the potential outcomes.
@@
In this work, we compare four different learning methods to estimate $\hat{\mu}$: an S-learner with random forests as the base learner (S-Learner (rf)), an S-learner with a feedforward multi-layer perceptron (MLP) without any debiasing as the base learner (S-Learner (mlp)) (Künzel et al., 2019), DRNet (Schwab et al., 2020), and VCNet (Nie et al., 2021). Further details on the hyperparameters can be found in E.
@@ 4.2 ILP for the dose-allocation problem
Deciding on the optimal treatment dose for each entity, represented by policies $\stackrel{{\scriptstyle B}}{{\Pi}}^{exp}$ and $\stackrel{{\scriptstyle B}}{{\Pi}}^{opt}$, is equivalent to solving Equations (9-10). In the case when only binary doses can be applied, i.e., $S\in\{0,1\}$, this problem is an instance of the knapsack problem (Dantzig, 1957). A sensible heuristic would be to rank allocations from high to low uplift until the budget capacity is reached. This approach mirrors the calculation of a traditional Qini curve in a binary treatment setting. In our more general setting, however, any dose $S\in[0,1]$ is permitted. To formalize this, we introduce two related functions:
$\stackrel{{\scriptstyle}}{{U}}:\{0,1\}^{(N\times\delta)}\times[-1,1]^{(N\times\delta)}\rightarrow\mathbb{R}\quad\text{and}\quad\stackrel{{\scriptstyle}}{{V}}:\{0,1\}^{(N\times\delta)}\times[-1,1]^{(N\times\delta)}\times\mathbb{R}^{N}\rightarrow\mathbb{R}.$
Both take as input a policy $\stackrel{{\scriptstyle B}}{{\Pi}}\in\{0,1\}^{(N\times\delta)}$ (expressed as a matrix of decision variables) and a matrix $\mathbf{T}\in[-1,1]^{(N\times\delta)}$ of CADE values (estimated $\hat{\tau}$ or ground-truth $\tau$) for the $N$ entities. The function $\stackrel{{\scriptstyle}}{{V}}$ additionally requires a benefit vector $\mathbf{b}\in\mathbb{R}^{N}$. Note that $\stackrel{{\scriptstyle}}{{U}}$ is a special case of $\stackrel{{\scriptstyle}}{{V}}$ in which the benefit vector $\mathbf{b}$ is set to all ones. For simplicity, we elaborate on $\stackrel{{\scriptstyle}}{{V}}$, with the understanding that this case also encompasses $\stackrel{{\scriptstyle}}{{U}}$.
@@
To solve Equations (910), we propose the following ILP (Equations 11-18), where $V_{i}(\pi)$ in the objective is replaced by $V_{i}^{exp}$ when finding $\stackrel{{\scriptstyle B}}{{\Pi}}^{exp}$ and by $V_{i}^{opt}$ when finding $\stackrel{{\scriptstyle B}}{{\Pi}}^{opt}$. Additionally, we compare the ILP to a greedy ranking heuristic (see Section 5.3.1) Although numerous business requirements exist, we focus on fairness as the main example of constraints throughout this work. Without loss of generalization, we consider one protected binary sensitive attribute $A$ (which is also included as an input feature to train the predictive model), where $N_{\mathrm{A=0}}$ is the index set of entities where the protected attribute $A=0$, and $N_{\mathrm{A=1}}$ represents those with $A=1$. These two sets are mutually exclusive and collectively exhaustive, so that $|N_{\mathrm{A=0}}|+|N_{\mathrm{A=1}}|=|N|$. To address fairness in allocation and outcomes, we introduce two slack parameters: $\epsilon_{DT}\in[0,1]$ for allocation fairness and $\epsilon_{DO}\in[0,1]$ for outcome fairness. A slack parameter value of $\epsilon=0$ guarantees perfect fairness between the groups $N_{\mathrm{A=0}}$ and $N_{\mathrm{A=1}}$, while parameter $\epsilon=1$ removes this fairness constraint.
max $\sum_{i=1}^{N}V_{i}(\pi)$ (11)
s.t. $\sum_{i=1}^{N}\Psi_{i}(\pi)\leq B$ (Budget constraint) (12)
$\frac{1}{|N_{\mathrm{A=0}}|}\sum_{i\in N_{\mathrm{A=0}}}\sum_{s\in D}\pi_{s}(X_{i})\cdot s\geq(1-\epsilon_{DT})\cdot\frac{1}{|N_{\mathrm{A=1}}|}\sum_{i\in N_{\mathrm{A=1}}}\sum_{s\in D}\pi_{s}(X_{i})\cdot s$ (Alloc. fairness 1) (13)
$\frac{1}{|N_{\mathrm{A=0}}|}\sum_{i\in N_{\mathrm{A=0}}}\sum_{s\in D}\pi_{s}(X_{i})\cdot s\leq(1+\epsilon_{DT})\cdot\frac{1}{|N_{\mathrm{A=1}}|}\sum_{i\in N_{\mathrm{A=1}}}\sum_{s\in D}\pi_{s}(X_{i})\cdot s$ (Alloc. fairness 2) (14)
$\frac{1}{|N_{\mathrm{A=0}}|}\sum_{i\in N_{\mathrm{A=0}}}\sum_{s\in D}\pi_{s}(X_{i})\cdot\tau_{s}\geq(1-\epsilon_{DO})\cdot\frac{1}{|N_{\mathrm{A=1}}|}\sum_{i\in N_{\mathrm{A=1}}}\sum_{s\in D}\pi_{s}(X_{i})\cdot\tau_{s}$ (Outc. fairness 1) (15)
$\frac{1}{|N_{\mathrm{A=0}}|}\sum_{i\in N_{\mathrm{A=0}}}\sum_{s\in D}\pi_{s}(X_{i})\cdot\tau_{s}\leq(1+\epsilon_{DO})\cdot\frac{1}{|N_{\mathrm{A=1}}|}\sum_{i\in N_{\mathrm{A=1}}}\sum_{s\in D}\pi_{s}(X_{i})\cdot\tau_{s}$ (Outc. fairness 2) (16)
$\sum_{s\in D}\pi_{s}(X_{i})=1,\forall i\in\{1,\ldots,N\}$ (Exactly one dose) (17)
$\pi_{s}(X_{i})\in\{0,1\},\forall s\in D,\forall i\in\{1,\ldots,N\}$ (Binary decisions) (18)
@@
Analogous to $V_{i}(\pi)^{exp}$, $V_{i}(\pi)^{presc}$, $V_{i}(\pi)^{opt}$ (Equations (6-8), which are defined on instance-level), and Equations (9-10) in Section 3, we consider three versions of $\stackrel{{\scriptstyle}}{{V}}$ each differing in their dependence on ground-truth or estimated CADEs:
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\stackrel{{\scriptstyle}}{{V}}^{exp}=\sum_{i=1}^{N}\sum_{d=1}^{\delta}\bigl{(}\stackrel{{\scriptstyle B}}{{\Pi}}^{exp}_{i,d}\cdot\hat{T}_{i,d}\cdot b_{i}\bigr{)},}$ (19)
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\stackrel{{\scriptstyle}}{{V}}^{presc}=\sum_{i=1}^{N}\sum_{d=1}^{\delta}\bigl{(}\stackrel{{\scriptstyle B}}{{\Pi}}^{exp}_{i,d}\cdot{T}_{i,d}\cdot b_{i}\bigr{)},}$ (20)
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\stackrel{{\scriptstyle}}{{V}}^{opt}=\sum_{i=1}^{N}\sum_{d=1}^{\delta}\bigl{(}\stackrel{{\scriptstyle B}}{{\Pi}}^{opt}_{i,d}\cdot{T}_{i,d}\cdot b_{i}\bigr{)}.}$ (21)
@@ 5.1 Data
Our experiments use a semi-synthetic approach based on a real dataset about infant health and development programs (IHDP) (Brooks-Gunn et al., 1992). This dataset originates from a randomized controlled trial and is frequently used as a benchmark for binary CATE estimation methods (Shalit et al., 2017; Guo et al., 2020). Since we focus on continuous treatments, we adopt a semi-synthetic approach where both doses and outcomes are artificially generated. For this purpose, we follow the established literature (Nie et al., 2021). Detailed information about the original dataset and the semi-synthetic data generation process is provided in C.